Please show step by step solution

Question 24 A 60 kg man and a 40 kg woman stand side by side at the same end of a 100 kg boat, and are ready to dive with velocities relative to the boat of 8 m/s and 5 m/s, respectively Determine the velocity of the boat after they have both dived if the woman dives first.
Answer #1

let the speed of the boat be V m/s

by the conservation of momentum we have

m1u1 + m2u2 = Mv

60*8 + 40*5 = 100*v

v = 6.8 m/s

so the boat will move with 6.8 m/s